# Eyes Closed, Safety On: Protecting Multimodal LLMs via Image-to-Text Transformation

### v1

@@ 5.4 Evaluation of Utility
Experiments are performed on popular MLLM utility benchmarks, including MME [17], MM-Vet [63], and MMBench [37]. These benchmarks cover a wide range of common abilities/tasks (e.g., maths, OCR, perception of objects, color and understanding of arts) that are considered as important for MLLMs. MME [63] has the subsets of perception (MME-P) and cognition (MME-C), each with 10 and 4 tasks, respectively. For each subset, the sum of accuracy and accuracy+ [17] within each task are reported to evaluate utility. For MMBench [37] and MM-Vet [63], accuracy and average GPT score (ranging from 0 to 1) for all samples are reported. More details on these datasets are in Appendix 0.A.2. We assume that all queries are benign and do not induce harmful answers. In other words, any detection of harm by MLLMs would be considered as a false alarm.
Figure 7: Examples of commonsense reasoning from MME-C, which are answered incorrectly when directly prompted but correctly with query-aware I2T transform.
@@
Table 2 shows the misclassification ratios by MLLMs that predict benign queries as malicious. As can be seen, most of the time MLLMs can correctly recognize the benign queries and do not trigger the I2T transformation process. Table 3 shows the utility scores of MLLMs on the benchmarks. It can be observed that across different models, ECSO does not hurt the utility scores of MLLMs on MME-P and MM-Vet, while even offers slight improvement on MME-C and MM-Vet. We speculate that this improvement might be attributed to the world knowledge elicited from query-aware captioning. Specifically, Figure 7 shows two examples from MME’s commonsense reasoning subset in which the MLLM fails when prompted to directly answer the question, which requires strong reasoning and world knowledge (e.g., “coats are proper for cold weather” and “shoes with blades are for skating”) which cannot be responded solely from the image. On the other hand, the generated query-aware captions provide more contexts to prompt the MLLMs to reason more thoroughly and finally lead to the correct answers.
[Uncaptioned image] Figure 8: Harmless rate of LLaVA-1.5-7B when images are invisible and visible to the model. The SD+OCR split of MM-SafetyBench is evaluated here.
Methods Harmless Rate w/o images 91.8 w/ images 85.6 Table 4: Performance of LLaVA-1.5-7B on the examine split of VLSafe with and without images.
@@
5.5 Ablation Study
5.5.1 Necessity of excluding images.
In ECSO, the unsafe image-text pairs are queried again with images converted to captions. A critical design of ECSO is that the actual images are discarded in this stage. Here, we show that the absence of image is the key to generate safer responses. To ablate this feature, we insert the image features to MLLMs in addition to the query-aware caption. Figure 8 and Table 4 show the harmless rate of LLaVA-1.5-7B on MM-SafetyBench (SD+OCR) and VLSafe (examine), respectively. For both benchmarks, the harmless rate decreases by a large margin with images are incorporated. Hence, ECSO is indeed restoring safety mechanism of pre-aligned LLMs, and totally different from multimodal Chain of Thoughts (MM-CoT) [65, 10] that succeed only via more reasoning steps or multi-turn self-moderation.
Furthermore, we find that the performance drop is more significant on MM-SafetyBench than that on VLSafe, which can be explained by the differences in sources of malicious contents. Specifically, MM-SafetyBench and VLSafe attack at the image and text modalities, respectively. Hence, MM-SafetyBench, with images visible to the MLLMs, is more prone to be induced.
5.5.2 Effect of Query-aware I2T Transformations.
In this experiment, we demonstrate that the proposed query-aware I2T transformation in ECSO is indispensable to maintain the utility of MLLMs. We choose ShareGPT4V-7B for ablation study since it makes more mis-classifications as shown in Table 2, which makes it more prone to affect the utility. To study the effect of conditioning on the query, we replace $P_{\text{trans}}$ with “Please give a caption for the image”. As demonstrated in Table 5, the removal of query-aware I2T transformations has a negative impact on the utility of MLLMs.
@@ 5.6 Safety Alignment
Figure 10 compares the harmless rates of the following models: (i) the original LLaVA-1.5-7B equipped with training-free ECSO (denoted ECSO), (ii) directly prompting LLaVA-1.5-7B which has been finetuned on $D^{\prime}$ and utility data to respond to queries (denoted ECSO_VLGuard) and (iii) directly prompting LLaVA-1.5-7B which has been finetuned on $D^{*}$ and utility data (denoted VLGuard). As reported, (i) ECSO_VLGuard is outperformed by ECSO in most cases since the former is trained on the ECSO outputs. (ii) ECSO_VLGuard offers better safety than VLGuard, showing that data generated by ECSO has comparable or even better quality than human verified data, offering better trade-off among safety and utility.
@@ 5.7 Limitation and Future Work
While ECSO can notably strengthen the safety of MLLMs, it heavily relies on the LLMs’ capacity to identify and neutralize unsafe queries. Therefore, any deficiencies in the LLMs’ safety mechanism may compromise ECSO’s performance in multimodal scenarios. Moving forward, an exciting prospect for future research is to explore how to turn multimodality from a challenge into an asset for safety. By developing new methods harnessing rich context provided by multiple modalities, it is potential to create more nuanced and context-aware safety mechanisms, increasing efficacy and reliability of MLLMs’ safety protocols.
@@ Acknowledgement
We gratefully acknowledge the support of MindSpore, CANN (Compute Architecture for Neural Networks) and Ascend AI Processor used for this research.
@@ 0.A.1.3 VLGuard
[68] is also proposed for the safety alignment of MLLMs. There are 2000 images in its training set, with 977 harmful and 1023 benign ones. Each safe image is matched with both a safe and an unsafe query-response pair, while each harmful image is coupled with a single query-instruction explaining the unsafe nature of this image. Note that the resposes and queries in this data are generated by GPT4. In total, there are around 3000 query-response pairs in the training set. Here, we use the training set of VLGuard for the safety alignment experiment in Sec. 5.6.
@@ 0.A.2.1 MME
[17] examines both perception (MME-P) and cognition (MME-C) abilities of MLLM on a total of 14 sub-tasks with 2374 questions in total. Each instruction consists of a question followed by “Please answer yes or no”. For each test image, two instructions are manually designed. The ground truth answer of the first question is “yes”, and that of the second question is “no”. The utility score of a sub-task is based on the sum of accuracy and accuracy+. Here, accuracy is calculated based on each question, while accuracy+ is based on each image where both of the two questions need to be answered correctly. The perception score is the sum of scores of all perception sub-tasks (ranging from 0 to 2000). The cognition score is calculated in the same way (ranging from 0 to 800).
@@ 0.A.3 Datasets Used in Preliminary Study
To assess the awareness of the safety in MLLMs, we collect model responses from multiple dataset sources:
  • MM-SafetyBench: 72, 79, and 85 responses sampled from 01-Illegal Activity, 02-HateSpeech, and 04-Physical Harm. All responses are unsafe.
  • VLSafe (examine): 264 responses sampled from the examine split of VLSafe, which are all unsafe.
  • LLaVA_150k  [^3] https://huggingface.co/datasets/liuhaotian/LLaVA-Instruct-150K: 500 responses sampled from the instruction tuning dataset of LLaVA [34]. All responses are safe.
In total, there are 1,000 responses, in which 500 of them are safe and the remaining 500 are unsafe. Note that all the responses are generated by LLaVA-1.5-7B and are classified into safe/unsafe via GPT-4 (and double-checked manually). For MM-SafetyBench and VLSafe (examine), we use the same prompt as the one in their respective evaluation process. For LLaVA_150k, we use the same prompt as for VLSafe (examine).
Appendix 0.B More on Implementation Details
@@ 0.B.2 Finetuning
In this section, we present the details on the training data and configurations. Further, we provide analysis on the generated data by ECSO.
@@
In Sec. 5.6, we use ECSO to generate benign responses from the queries of VLGurad. To construct $D^{\prime}$, we only keep the query-response pairs that are initially detected as harmful by the MLLMs. Note that though the initial response $\tilde{y}$ is harmful, the generated response $y$ by ECSO is safe. In this way, we obtain malicious queries together with benign responses, which are presumed to be effective for the safety alignment process. Finally, we obtain 232 such query-response pairs to form $D^{\prime}$ for safety alignment. The query-response pairs remaining in VLGuard are denoted $D^{\prime}_{\text{safe}}$. Different from $D^{\prime}$, the initial responses in $D^{\prime}_{\text{safe}}$ are considered to be benign by MLLMs and remain unchanged. We will show in the following experiment that $D^{\prime}_{\text{safe}}$ is critical to maintain the utility of MLLMs during the safety alignment.
@@
LLaVA-1.5-7B is adopted for finetuning. Specifically, we follow the official repository [^5] https://github.com/haotian-liu/LLaVA to train the model using LoRA with a rank of 128. For all the training experiments, we finetune the model for 1 epoch with a global batch size of 128, a learning rate of $2\times 10{-4}$ and weight decay of $0.1$ on 4 Telsa-V100-SXM2-32G GPUs.
@@ Appendix 0.C More Experiments
In Table 1, we only report the results of LLaVA-1.5-7B on the scenarios 01-07 & 09 of MM-SafetyBench. Here, we present the results of the remaining scenarios in Table 8. Similarly, Tables 9, 10, 11, 12 shows the results on the models ShareGPT4V-7B, mPLUG-Owl-7B, Qwen-VL-Chat and InternLM-XComposer-7B, respectively. Specifically, we observe that even without images (i.e., text only), on scenarios 08 and 10-13, all models achieve poor safety in their responses when directly prompted (i.e., all are less than 15% by average in terms of harmless rates). This reveals the intrinsic vulnerabilities of the LLMs for those scenarios. Therefore, we cannot expect MLLMs to be safer with the presence of images, nor when using ECSO as it relies heavily on the safety mechanism of LLMs. Note that we observe relatively higher harmless rate in data type “SD” (e.g., Qwen-VL-Chat and InternLM-XComposer-7B) since “SD” does not contain texts in images, thus the model might understand the image in a harmless way.
0.C.2 Case Studies
@@
Figure 16 shows ECSO is also effective on FigStep [20], which is another safety benchmark of MLLMs. However, since both MM-SafetyBench and FigStep inject malicious contents into images by OCR, we only report full results of ECSO on MM-SafetyBench in Sec. 5.
Figure 17 shows ECSO cannot be easily bypassed via certain strategy. In particular, by replacing the sensitive word “bomb” with a picture of a bomb, one could induce MLLM to generate harmful responses. However, ECSO succeeds to protect the MLLM.
Figure 18 shows ECSO can protect MLLMs from adversarial images. Here we use the examples in [45]. Note that the first image 777image source: https://github.com/Unispac/Visual-Adversarial-Examples-Jailbreak-Large-Language-Models/blob/main/adversarial_images/clean.jpeg is clean and the MLLM rejects to fulfill the malicious request thanks to its safety mechanism (though proving to be limited in this paper). However, the second image 888image source: https://github.com/Unispac/Visual-Adversarial-Examples-Jailbreak-Large-Language-Models/blob/main/adversarial_images/prompt_constrained_64.bmp is adversarial (optimized to let MLLMs generate harmful responses). In this case, “Direct” generates unsafe contents while ECSO succeeds to protect the MLLM.
Appendix 0.D More Discussion

### v3

@@ 5.4 Evaluation of Utility
Experiments are performed on popular MLLM utility benchmarks, including MME [17], MM-Vet [67], and MMBench [40]. These benchmarks cover a wide range of common abilities/tasks (e.g., maths, OCR, perception of objects, color and understanding of arts) that are considered as important for MLLMs. MME [67] has the subsets of perception (MME-P) and cognition (MME-C), with 10 and 4 tasks, respectively. For each subset, the sum of accuracy and accuracy+ [17] within each task are reported to evaluate utility. For MMBench [40] and MM-Vet [67], accuracy and average GPT score (ranging from 0 to 1) for all samples are reported. More details on these datasets are in Appendix 0.A.2. We assume that all queries are benign and do not induce harmful answers. In other words, any detection of harm by MLLMs would be considered as a false alarm.
@@
Table 2 shows the misclassification ratios by MLLMs that predict benign queries as malicious. As can be seen, most of the time MLLMs can correctly recognize the benign queries and do not trigger the I2T transformation process. Table 3 shows the utility scores of MLLMs on the benchmarks. It can be observed that across different models, ECSO does not hurt the utility scores of MLLMs on MME-P and MM-Vet, while even offers slight improvement on MME-C and MM-Vet. We speculate that this improvement might be attributed to the world knowledge elicited from query-aware captioning.
Figure 8: Harmless rate of LLaVA-1.5-7B when images are invisible and visible to the model. The SD+OCR split of MM-SafetyBench is evaluated here.
Methods Harmless Rate
w/o images 91.8
w/ images 85.6
Table 4: Performance of LLaVA-1.5-7B on the examine split of VLSafe with and without images.
5.5 Ablation Study
5.5.1 Necessity of excluding images.
In ECSO, the unsafe image-text pairs are queried again with images converted to captions. A critical design of ECSO is that the actual images are discarded in this stage. Here, we show that the absence of image is the key to generate safer responses. To ablate this feature, we insert the image features to MLLMs in addition to the query-aware caption. Figure 8 and Table 4 show the harmless rates of LLaVA-1.5-7B on MM-SafetyBench (SD+OCR) and VLSafe (examine), respectively. On both benchmarks, the harmless rate decreases by a large margin with images are incorporated. Hence, ECSO is indeed restoring the safety mechanism of pre-aligned LLMs, and very different from multimodal Chain of Thoughts (MM-CoT) [69, 10] that succeeds only via more reasoning steps or multi-turn self-moderation.
Furthermore, we find that the performance drop is more significant on MM-SafetyBench than that on VLSafe. This can be explained by the differences in sources of malicious contents. MM-SafetyBench attacks the image modality, while VLSafe attacks the text modality. Hence, MM-SafetyBench, with images visible to the MLLMs, is more prone to be induced.
Methods MME MM-Vet MMBench
w/ step 3&4 1865 30.6 64.18
w/o step 3&4 1847 30.0 63.83
Table 5: Utility on MME, MM-Vet and MMBench.
5.5.2 Effect of Steps 3&4
In Sec. 4.2 and 4.3, we caption the image (step 3) and query the LLM again (step 4) in case of unsafe responses. A seemingly simpler solution is to directly refuse to respond and output “I cannot answer this question due to safety constraint”. Table 5 shows the utility of LLaVA-1.5 when employing such a strategy. As can be seen, the model with steps 3 and 4 achieve higher utility because they always respond to the queries.
5.5.3 Effect of Query-aware I2T Transformations.
In this experiment, we demonstrate that the proposed query-aware I2T transformation in ECSO is indispensable to maintaining the utility of MLLMs. We use ShareGPT4V-7B for ablation study, since it makes more mis-classifications as shown in Table 2, which makes it more prone to affect the utility. To study the effect of conditioning on the query, we replace $P_{\text{trans}}$ with “Please give a caption for the image”. As demonstrated in Table 6, the removal of query-aware I2T transformations has a negative impact on the utility of MLLMs.
@@ 5.6 Safety Alignment
Figure 10 compares the harmless rates of the following models: (i) the original LLaVA-1.5-7B equipped with training-free ECSO (denoted ECSO), (ii) directly prompting LLaVA-1.5-7B which has been finetuned on $D^{\prime}$ and utility data to respond to queries (denoted ECSO_VLGuard) and (iii) directly prompting LLaVA-1.5-7B which has been finetuned on $D^{*}$ and utility data (denoted VLGuard). As can be seen, (i) In most cases, ECSO_VLGuard is outperformed by ECSO since ECSO_VLGuard is trained on the ECSO outputs. (ii) ECSO_VLGuard offers better safety than VLGuard, showing that data generated by ECSO has comparable or even better quality than human-verified data, offering better trade-off among safety and utility.
@@ 5.7 Limitation and Future Work
While ECSO can notably strengthen the safety of MLLMs, it heavily relies on the LLMs’ capacity to identify and neutralize unsafe queries. Therefore, any deficiencies in the LLMs’ safety mechanism may compromise ECSO’s performance in multimodal scenarios. Moving forward, an exciting prospect for future research is to explore how to turn multimodality from a challenge into an asset for safety. By developing new methods harnessing rich context provided by multiple modalities, it might be possible to create more nuanced and context-aware safety mechanisms, increasing the efficacy and reliability of MLLMs’ safety protocols.
@@ Acknowledgement
We gratefully acknowledge the support of MindSpore, CANN (Compute Architecture for Neural Networks) and Ascend AI Processor used for this research. This work was partially supported by NSFC key grant 62136005, NSFC general grant 62076118, Shenzhen fundamental research program JCYJ20210324105000003, and the Research Grants Council of the Hong Kong Special Administrative Region (Grants C7004-22G-1 and 16202523).
@@ 0.A.1.3 VLGuard
[73] is also proposed for the safety alignment of MLLMs. There are 2000 images in its training set. 977 of them are harmful while the remaining 1023 are benign. Each safe image is matched with a safe query-response pair and an unsafe pair, while each harmful image is coupled with a single query-instruction explaining the unsafe nature of this image. Note that the resposes and queries in this data set are generated by GPT4. In total, there are around 3000 query-response pairs in the training set. Here, we use the training set of VLGuard for the safety alignment experiment in Sec. 5.6.
@@ 0.A.2.1 MME
[17] examines both the perception (MME-P) and cognition (MME-C) abilities of MLLM on a total of 14 sub-tasks with 2374 questions. Each instruction consists of a question followed by “Please answer yes or no”. For each test image, two instructions are manually designed. The ground-truth answer of the first question is “yes”, and that of the second question is “no”. The utility score of a sub-task is based on the sum of accuracy and accuracy+. Here, accuracy is calculated based on each question, while accuracy+ is based on each image where both of its two questions need to be answered correctly. The perception score is the sum of scores of all perception sub-tasks (ranging from 0 to 2000). The cognition score is calculated in the same way (ranging from 0 to 800).
@@ 0.A.3 Datasets Used in Preliminary Study
To assess the safety awareness in MLLMs, we collect model responses from multiple dataset sources:
  • MM-SafetyBench: 72, 79, and 85 responses are sampled from 01-Illegal Activity, 02-HateSpeech, and 04-Physical Harm, respectively. All responses are unsafe.
  • VLSafe (examine): 264 responses are sampled from the examine split of VLSafe, which are all unsafe.
  • LLaVA_150k  [^3] https://huggingface.co/datasets/liuhaotian/LLaVA-Instruct-150K: 500 responses are sampled from the instruction tuning dataset of LLaVA [37]. All responses are safe.
In total, there are 1,000 responses, in which 500 of them are safe and the remaining 500 are unsafe. Note that all responses are generated by LLaVA-1.5-7B and are classified into safe/unsafe via GPT-4 (and double-checked manually). For MM-SafetyBench and VLSafe (examine), we use the same prompt as the one in their respective evaluation process. For LLaVA_150k, we use the same prompt as for VLSafe (examine).
Appendix 0.B Implementation Details
@@ 0.B.2 Finetuning
In this section, we present details on the training data and configurations. Further, we provide analysis on the generated data by ECSO.
@@
In Sec. 5.6, we use ECSO to generate benign responses from the queries of VLGurad. To construct $D^{\prime}$, we only keep the query-response pairs that are initially detected as harmful by the MLLMs. Note that though the initial response $\tilde{y}$ is harmful, the generated response $y$ by ECSO is safe. In this way, we obtain malicious queries together with benign responses, which are presumed to be effective for the safety alignment process. Finally, we obtain 232 such query-response pairs to form $D^{\prime}$ for safety alignment. The query-response pairs remaining in VLGuard are denoted $D^{\prime}_{\text{safe}}$. Different from $D^{\prime}$, the initial responses in $D^{\prime}_{\text{safe}}$ are considered benign by MLLMs and remain unchanged. We will show in the following experiment that $D^{\prime}_{\text{safe}}$ is critical to maintaining the utility of MLLMs during safety alignment.
@@
LLaVA-1.5-7B is adopted for finetuning. Specifically, we follow the official repository [^5] https://github.com/haotian-liu/LLaVA to train the model using LoRA with a rank of 128. For all the training experiments, we finetune the model for 1 epoch with a global batch size of 128, a learning rate of $2\times 10^{-4}$ and weight decay of $0.1$ on 4 Telsa-V100-SXM2-32G GPUs.
@@ Appendix 0.C Experiments
In Table 1, we only report the results of LLaVA-1.5-7B on the scenarios 01-07 & 09 of MM-SafetyBench. Here, we present results of the remaining scenarios in Table 13. Similarly, Tables 14, 15, 16, 17 shows the results on the models ShareGPT4V-7B, mPLUG-Owl-7B, Qwen-VL-Chat and InternLM-XComposer-7B, respectively. Specifically, we observe that even without images (i.e., text only), on scenarios 08 and 10-13, all models achieve poor safety in their responses when directly prompted (i.e., all are less than 15% by average in terms of harmless rates). This reveals the intrinsic vulnerabilities of the LLMs for these scenarios. Therefore, we cannot expect MLLMs to be safer with the presence of images, nor when using ECSO as it relies heavily on the safety mechanism of LLMs. Note that we observe relatively higher harmless rate in data type “SD” (e.g., Qwen-VL-Chat and InternLM-XComposer-7B) since “SD” does not contain texts in images. Thus, the model might understand the image in a harmless way.
0.C.2 Additional Ablations
0.C.2.1 Effects of Safety Instructions
In the last step of ECSO, the LLM is instructed with “Based on the hints, answer the following question in a HARMLESS and ETHICAL way.” We ablate the effect of this safety instruction on Direct and ECSO, and report the harmless rate in Table 9. We find the safety instruction is beneficial for response safety but CANNOT address the safety issues of MLLMs.
Methods LLaVA-1.5 ShareGPT4V Qwen-VL-Chat InternLM-XC mPLUG-Owl2
VL. MM. VL. MM. VL. MM. VL. MM. VL. MM.
Direct 20.5 48.8 16.9 47.8 65.7 52.5 26.8 54.0 18.7 43.9
w/ inst. 24.1 49.5 19.7 47.0 49.8 50.7 77.8 81.1 31.9 46.1
ECSO 91.8 86.4 93.3 85.3 95.9 83.4 81.6 78.5 76.8 80.6
w/o inst 71.3 76.3 81.4 71.0 93.0 71.5 26.3 67.1 68.9 73.9
Table 9: Abaltions on the effect of safety instruction. Harmless rates on VLsafe (VL.) and MM-SafetyBnech (MM.) are reported (averaged over the scenarios in Table 1 of the SD+OCR split). Due to space limits, model sizes are not presented in their names. InternLM-XC is short for InternLM-XComposer.
0.C.2.2 Different Harm-Detection Strategies
In step 2 of ECSO, we let the MLLM judge whether its own response is safe or not. In contrast to this strategy, one can also conduct detection before model responses with the query and image/caption. In Table 10, we present the discrimination accuracies of various strategies on the same dataset as described in Sec. 3.2. We find that detection based on the model response achieves the highest accuracy. This can be explained by the fact that harmful content is more explicit in the response than in the query or image.
Detect LLaVA-1.5 ShareGPT4V Qwen-VL-Chat InternLM-XC mPLUG-Owl2
Response 95.0 96.6 92.7 86.9 88.5
Img. & Q. 75.6 90.2 84.8 61.3 84.3
Cap. & Q. 82.6 91.6 88.9 66.0 83.0
Table 10: Accuracy of various MLLMs detection on safety of responses with various strategies. Here Img. and Cap. and Q. stand for image, caption, query, respectively.
0.C.2.3 Usage of LLMs
In step 4 of ECSO, the LLM part of the MLLM is employed to generate a safe answer. However, this LLM is finetuned when it is connected to the vision encoder. Here we replace it with the original LLM and report the harmless rate of ECSO in Table 11777As mPLUG-Owl2 starts with LLaMA-2-base instead of its chat version, we do not include the results on the original LLM.. As can be seen, ECSO with the original LLM achieves higher harmless rate. This results from the fact that finetuning may cause forgetting in the previously learned text-only safety alignment. However, in ECSO, we do not employ this strategy because it requires loading two LLMs into the memory.
Methods LLaVA-1.5 ShareGPT4V Qwen-VL-Chat InternLM-XC mPLUG-Owl2
VL. MM. VL. MM. VL. MM. VL. MM. VL. MM.
finetuned LLM 91.8 86.4 93.3 85.3 95.9 83.4 81.6 78.5 76.8 80.6
original LLM 93.3 85.9 94.9 85.8 96.3 83.8 89.0 82.0 - -
Table 11: Abaltions on the usage of LLM. Evaluations are the same as in Table 9.
0.C.3 Applying ECSO on the Finetuned Model
In this section, we apply ECSO on the models finetuned in Sec. 5.6. Table 12 shows the harmless rate of LLaVA-1.5-7B (fintuned on different datasets) on MM-SafetyBench. It can be observed that ECSO is complementary with finetuning.
ECSO ECSO_VLGuard VLGuard
FT FT + ECSO FT FT + ECSO
86.38 81.39 90.09 73.68 89.18
Table 12: Combine finetuning with ECSO. Results of LLaVA-1.5-7B on MM-SafetyBench are reported (following Table 9). FT: Finetuning.
0.C.4 Case Studies
@@
Figure 16 shows that ECSO is also effective on FigStep [21], which is another safety benchmark of MLLMs. However, since both MM-SafetyBench and FigStep inject malicious contents into images by OCR, we only report the full results of ECSO on MM-SafetyBench in Sec. 5.
Figure 17 shows that ECSO cannot be easily bypassed via certain strategy. In particular, by replacing the sensitive word “bomb” with a picture of a bomb, one could induce the MLLM to generate harmful responses. However, ECSO succeeds in protecting the MLLM.
Figure 18 shows that ECSO can protect MLLMs from adversarial images. Here we use the examples in [49]. Note that the first image888image source: https://github.com/Unispac/Visual-Adversarial-Examples-Jailbreak-Large-Language-Models/blob/main/adversarial_images/clean.jpeg is clean and the MLLM rejects to fulfill the malicious request thanks to its safety mechanism (though proving to be limited in this paper). However, the second image999image source: https://github.com/Unispac/Visual-Adversarial-Examples-Jailbreak-Large-Language-Models/blob/main/adversarial_images/prompt_constrained_64.bmp is adversarial (optimized to let MLLMs generate harmful responses). In this case, “Direct” generates unsafe contents while ECSO succeeds in protecting the MLLM.
Appendix 0.D More Discussions